OpenStudy (anonymous):

Solve 1.8x + 2.5 = 3.6

OpenStudy (anonymous):

1.8x=1.1 x = (1.1/1.8)

OpenStudy (anonymous):

the aim here is to 'leave x alone'

OpenStudy (anonymous):

that is why we have to do some arithmetic operations(know some rules etc): 1.8x=3.6-2.5 1.8x=1.1 x=1.1/1.8 x=11/18 opposite side means reverse(operation),ex:if it was plus in left side it should go as minus to right side

OpenStudy (anonymous):

11/18